# You Only Fine-tune Once: Many-Shot In-Context Fine-Tuning for Large Language Model

### v1

@@ 4 Experiments
This section compares our proposed many-shot in-context fine-tuning (ManyICL) with baseline methods on ICL performance. The datasets, baselines, and evaluation metrics are listed below.
43 Datasets: The experimental setup follows the MetaICL experiment (Min2022metaicl, ) and utilizes a large collection of tasks taken from CROSSFIT (ye2021crossfit, ). A total of 43 unique datasets are used, covering five task categories: text classification (CLS), question answering (QA), natural language inference (NLI), math (MATH), and summarization (SUM). Detailed descriptions of datasets split are in Appendix Section A.1 and Table 6, and the in-context example preprocessing can be found in Appendix Section A.2. The evaluation metrics are in Appendix Section A.3.
Baselines: The baseline methods for comparison are summarized in Table 3. Many-shot and few-shot fine-tuning include two variants: mask all targets and mask last target. Additionally, the many-shot fine-tuning setting includes a standard autoregressive variant that computes loss over all tokens. The training datasets for zero/few/many-shot are held-in datasets  [^1] The full datasets are listed in Appendix Section A. In contrast, task-level fine-tuning is performed on held-out datasets (14 separate models, one per task), which serve as an upper bound performance, since they are trained directly on the held-out tasks. We provide a definition for held-in and held-out data split. Held-in: 29 training datasets for zero/few/many-shot fine-tuning, these are listed in the training column of Table 6. Held-out: 14 datasets for evaluation across all methods. The full datasets are listed in the test column of Table 6.
@@
Table 4: Many-shot fine-tuning outperforms the zero/few-shot fine-tuning approaches. Metrics are reported as both the best accuracy and the accuracy achieved with the maximum number of in-context samples. Bold highlights the best performance across different meta-learning approaches (zero-, few-, and many-shot). The last row shows the performance of the dedicated fine-tuned model for each task. For NLI and QA, we report the average performance within the domain, and the concrete tasks in evaluation are listed in Table6.
CLS NLI QA MATH SUM
Clinc Banking77
@@
0.978 0.912 0.875 0.852 0.516 0.330
Zero-shot fine-tuning follows the standard multi-task learning approach. In all experiments, we use Mistral-v0.3 with 32K token length (jiang2023mistral, ) and use LoRA to fine-tune the model (hu2021lora, ). These are implemented with PyTorch and Transformers library. The LoRA parameters are listed below. The evaluation metric and in-context sample selection are provided in Appendix Section A.
Hyperparameters: We tuned hyperparameters (e.g., learning rate, weight decay, LoRA rank) for each baseline on a separate validation dataset. To ensure consistency across experiments, we fixed the batch size to 1 per rank due to memory constraints, used the same optimizer (default settings), and applied a cosine learning rate decay schedule for all models. All experiments are conducted on $8$ A100-80G GPUs, all with a per-device batch size of $1$ and a global batch size of $32$, using $4$ gradient accumulation steps. We use Adam optimizer with weight decay $10^{-5}$, $\beta_{1},\beta_{2}$ are $0.9$ and $0.99$, respectively.
@@ 4.1 Overall performance comparison
CLS: Figure 5(a-b) shows Classification (CLS) results on clinc (151 classes) and Banking77 (77 classes). While few-shot fine-tuning can outperform many-shot with less than 200 examples (consistent with MetaICL Min2022metaicl ), many-shot consistently excels with more samples. It reaches 0.968 peak accuracy (1500 examples), outperforming few-shot methods and validating our strategy. Notably, with >1200 examples, many-shot models on held-out data near task-level fine-tuning performance.
NLI: Figure 5(c) presents the experimental results for NLI task. The model is evaluated using varying numbers of in-context examples, ranging from 0 to 500 shots. Many-shot fine-tuning with the mask all targets strategy achieves the highest accuracy of 0.859 with 150 shots, outperforming the few-shot and zero-shot fine-tuning strategies by $3\%$. This improvement further narrows the performance gap between in-context learning (ICL) and task-level fine-tuning.
QA: Figure 5(d) shows the experimental results for the QA task, evaluated with 0 to 750 in-context examples. Notably, many-shot fine-tuning using the mask-all targets strategy achieves the highest performance, reaching an accuracy of 0.835 with 200 shots, which is $2.5\%$ higher than the best results obtained from few-shot and zero-shot fine-tuning. These results highlight the effectiveness of many-shot fine-tuning for QA tasks.
MATH: The results for the GSM8K dataset are presented in Figure 5(e). The performance of various fine-tuning strategies are evaluated across a range of 0 to 120 shots. It is observed that many-shot fine-tuning consistently achieves the best performance across all shots. The accuracy at the maximum number of shots is $0.467$, which is $1\%$ higher than the results of the few-shot fine-tuning method and $3.7\%$ higher than no-fine-tuned model.
SUM: For the summarization tasks, the evaluation is performed on the Newsroom dataset, with the results presented in Figure 5(f). We evaluate the performance from 0 to 25 shots, since for summarization, each shot contains much more tokens than other tasks. Results show that many-shot fine-tuning outperforms both zero-shot and few-shot scenarios and the performance gap between ICL and task-level fine-tuning is reduced.
@@ 4.2 Long-context capability comparison
We investigated how different fine-tuning strategies affect LLMs’ general capabilities, focusing on long-context understanding. LLMs are known to suffer from catastrophic forgetting during fine-tuning (zhai2023investigating, ; luo2023empirical, ). To evaluate this part, we use the PG-19 validation set (rae2019compressive, ) and report the perplexity of each method on this dataset, as shown in Table 5. Perplexity score at various context length has been a standard evaluation for long-context capability evaluation.
@@ 4.3 Ablation study
For the NLI task (Figure 6(c)), excluding NLI training datasets has little impact on performance. The ablation model’s curve (black line) nearly overlaps with the original model’s, demonstrating that many-shot ICL capabilities learned from other tasks generalize well to NLI. In the QA setting (Figure 6 (d)), the model fine-tuned without QA data still consistently outperforms the base model, suggesting strong transfer. However, incorporating QA data during fine-tuning can yield additional performance gains. For the MATH task in Figure 6(e), a similar pattern emerges: many-shot fine-tuning without direct MATH supervision still improves performance over the base model. We attribute this to transfer from other CoT-style (Chain-of-Thought) reasoning tasks in the training set.
These findings suggest that many-shot fine-tuning enables robust transfer to CLS, QA, MATH, and NLI, even without task-specific training data. However, as shown in Figure 6 (f), summarization (SUM) shows a notable exception. Removing summarization training datasets substantially degrades performance on summarization tasks, making it close to the base model’s level. We hypothesize two contributing factors: (1) SUM differs fundamentally from CLS, QA, NLI, and MATH, making it harder to transfer from unrelated tasks; and (2) SFT may be insufficient for developing strong summarization capabilities, potentially requiring more specialized training strategies (liu2020learning, ).
@@ 5 Conclusion
In this paper, we introduce ManyICL, a novel meta-training framework that enhances many-shot ICL for LLMs. Our approach shows superior performance in both few-shot and many-shot scenarios across diverse downstream tasks. Extensive experiments demonstrate that ManyICL outperforms conventional zero-shot and few-shot fine-tuning methods while mitigating catastrophic forgetting. These findings highlight ManyICL’s potential to streamline LLM deployment by reducing reliance on task-specific fine-tuning.
@@
Acknowledgments
This research was supported by the Google Student Researcher Program. We are grateful for the valuable discussions and guidance provided by Yinpeng Chen and Jesper Andersen, and for the leadership support from Jay Yagnik and Prateek Jain.

### v2

@@ 4 Experiments
This section compares our proposed many-shot in-context fine-tuning (ManyICFT) with baseline methods on ICL performance. The datasets, baselines, and evaluation metrics are listed below.
43 Datasets: The experimental setup follows the MetaICL experiment [Min2022metaicl] and utilizes a large collection of tasks taken from CROSSFIT [ye2021crossfit]. A total of 43 unique datasets are used, covering five task categories: text classification (CLS), question answering (QA), natural language inference (NLI), multi-label classification (ML-CLS) and multilingual summarization (SUM). Detailed descriptions of datasets split are in Appendix Section A.1 and Table 6, and the in-context example preprocessing can be found in Appendix Section A.2. The evaluation metrics are in Appendix Section A.3.
Baselines: The baseline methods for comparison are summarized in Table 3. Many-shot and few-shot fine-tuning include two variants: mask all targets and mask last target. Additionally, the many-shot fine-tuning setting includes a standard autoregressive variant that computes loss over all tokens. The training datasets for zero/few/many-shot are held-in datasets  [^1] The full datasets are listed in Appendix Section A. In contrast, task-level fine-tuning is performed on held-out datasets (14 separate models, one per task), which serve as an upper bound performance, since they are trained directly on the held-out tasks. We provide a definition for held-in and held-out data split. Held-in: 30 training datasets for zero/few/many-shot fine-tuning, these are listed in the training column of Table 6. Held-out: 14 datasets for evaluation across all methods. The full datasets are listed in the test column of Table 6. For multilingual summarization task, we do fine-tuning on the high-resource language and eval on the low-resource datasets.
@@
Table 4: Many-shot fine-tuning outperforms the zero/few-shot fine-tuning approaches. Metrics are reported as both the best accuracy and the accuracy achieved with the maximum number of in-context samples (formatted in the table as Best Accuracy / Max Context Accuracy). Bold highlights the best performance across different meta-learning approaches (zero-, few-, and many-shot). The last row shows the performance of the dedicated fine-tuned model for each task. For NLI and QA, we report the average performance within the domain, and the concrete tasks in evaluation are listed in Table 6.
CLS NLI QA Multilingual SUM
Clinc Banking77 PT SR
@@
0.978 0.912 0.875 0.852 0.253 0.245
Zero-shot fine-tuning follows the standard multi-task learning approach. In all experiments, we use Mistral-v0.3 with 32K token length [jiang2023mistral] and use LoRA [hu2021lora] to finetune the model. These are implemented with PyTorch and Transformers library. The LoRA parameters are listed below. The evaluation metric and in-context sample selection are in Appendix Section A.
Hyperparameters: We tuned hyperparameters (e.g., learning rate, weight decay, LoRA rank) for each baseline on a separate validation dataset. To ensure consistency across experiments, we fixed the batch size to 1 per rank due to memory constraints, used the same optimizer (default settings), and applied a cosine learning rate decay schedule for all models. All experiments are conducted on $8$ A100-80G GPUs, all with a per-device batch size of $1$ and a global batch size of $32$, using $4$ gradient accumulation steps. We use Adam optimizer with weight decay $10^{-5}$, $\beta_{1},\beta_{2}$ is $0.9$ and $0.99$.
@@ 4.1 Overall performance comparison
CLS: Figure 5(a-b) shows Classification (CLS) results on clinc (151 classes) and Banking77 (77 classes). While few-shot fine-tuning can outperform many-shot with less than 200 examples (consistent with MetaICL Min2022metaicl), many-shot consistently excels with more samples. It reaches 0.968 peak accuracy (1500 examples), outperforming few-shot methods and validating our strategy. Notably, with >1200 examples, many-shot models on held-out data near task-level fine-tuning performance
NLI: Figure 5(c) presents the experimental results for NLI task. The model is evaluated using varying numbers of in-context examples, ranging from 0 to 500 shots. Many-shot fine-tuning with the mask all targets strategy achieves the highest accuracy of 0.868 with 150 shots, outperforming the few-shot and zero-shot fine-tuning strategies by $3\%$. This improvement further narrows the performance gap between in-context learning (ICL) and task-level fine-tuning.
QA: Figure 5(d) shows the experimental results for the QA task, evaluated with 0 to 750 in-context examples. Notably, many-shot fine-tuning using the mask-all targets strategy achieves the highest performance, reaching an accuracy of 0.845 with 200 shots, which is $2.5\%$ higher than the best results obtained from few-shot and zero-shot fine-tuning. These results highlight the effectiveness of many-shot fine-tuning for QA tasks.
Multilingual SUM: The results for the multilingual summarization dataset are presented in Figure 5(e) and (f). The performance of various fine-tuning strategies are evaluated across a range of 0 to 24 shots. It is observed that many-shot fine-tuning consistently achieves the best performance across all shots. The accuracy at the maximum number of shots is $0.467$, which is $1\%$ higher than the results of the few-shot fine-tuning method and $3.7\%$ higher than no-finetuned model.
@@ 4.2 Long-context capability comparison
We investigated how different fine-tuning strategies affect LLMs’ general capabilities, focusing on long-context understanding. LLMs are known to suffer from catastrophic forgetting during fine-tuning [zhai2023investigating, luo2023empirical]. To evaluate this part, we use the PG-19 validation set [rae2019compressive] and report the perplexity of each method on this dataset, as shown in Table 5. Perplexity score at various context length has been a standard evaluation for long-context capability evaluation.
@@ 4.3 Ablation study
For the NLI task (Figure 6(c)), excluding NLI training datasets has little impact on performance. The ablation model’s curve (black line) nearly overlaps with the original model’s, demonstrating that many-shot ICL capabilities learned from other tasks generalize well to NLI. In the QA setting (Figure 6 (d)), the model fine-tuned without QA data still consistently outperforms the base model, suggesting strong transfer. However, incorporating QA data during fine-tuning can yield additional performance gains. For the Multilingual SUM task in Figure 6(e) and (f), a similar pattern emerges: many-shot fine-tuning without direct MATH supervision still improves performance over the base model. We attribute this to transfer from other CoT-style (Chain-of-Thought) reasoning tasks in the training set.
These findings suggest that many-shot fine-tuning enables robust transfer to CLS, QA, Multilingual SUM, and NLI, even without task-specific training data.
@@ 5 Conclusion
In this paper, we introduce ManyICFT, a novel meta-training framework that enhances many-shot ICL for LLMs. Our approach shows superior performance in both few-shot and many-shot scenarios across diverse downstream tasks. Extensive experiments demonstrate that ManyICFT outperforms conventional zero-shot and few-shot fine-tuning methods while mitigating catastrophic forgetting. These findings highlight ManyICFT’s potential to streamline LLM deployment by reducing reliance on task-specific fine-tuning.
@@
Appendix A Datasets
A.1 Datasets split
The training and test datasets are listed in Table 6. We used 43 datasets in total, including 29 training datasets and 14 test datasets among $5$ tasks. There is no overlap between the training and test datasets. For the training datasets, we select the training samples from the training split, and for the test datasets, we select the test samples from the test split of each dataset.
For many-shot fine-tuning, a total of 20K training instances are used (4K per task), with each instance composed of the maximum number of in-context examples that can fit within the model’s input context length. For few-shot fine-tuning setting, 5-shot prompts are applied across all datasets. On average, each training instance contains 212 in-context examples. This results in approximately 40 times more sequences for few-shot fine-tuning with the mask-all-targets strategy and 200 times more for zero-shot and few/many-shot fine-tuning with the mask-last-target strategy. This sampling approach ensures all baseline methods receive an equivalent number of training samples, enabling fair comparison across zero-shot, few-shot, and many-shot fine-tuning. For each instance, in-context examples are randomly selected from the training split. Datasets for each task are partitioned into non-overlapping training and test sets, allowing for evaluation of held-out generalization performance across all methods
A.2 ICL examples
$\bullet$ In-context sample selection: For the classification task, we use stratified sampling from training data, which samples the same number of samples for each class to ensure the balance of classes. For other tasks, we do random sampling from the training data. In-context samples are not fixed for different samples.
$\bullet$ Prompt construction: We construct the ICL prompts by concatenating the maximum number of input-output pairs that can fit within the model’s context window. The following illustrates the prompt template used for the Clinc dataset with a 3-shot example. Prompts for other datasets follow a similar construction methodology.
Task: You will receive a text input. Your task is to classify the
intent of the input text. Here are some examples:
Text: Let me know how much money I will need to spend on paying bills
Intent: Bill balance
Text: Could you let me know what song this is
Intent: What song
Text: Should I take a different route to work
Intent: Traffic
Text: {query}
Intent:
A.3 Evaluation metrics
For classification, QA, NLI and paraphrase detection tasks, we report the accuracy, and for summarization task, we report the ROUGE-L between prediction and ground-truth, which measures the longest overlap ratio between two texts.
Table 6: Datasets configurations. Training and test datasets for experiments.
Training (29 datasets)
Test (14 datasets)
CLS
Training: 10 datasets
Test: 2 datasets
hate_speech_offensive [de2018hate]
google_wellformed_query [faruqui2018identifying]
tweet_eval-emoji [barbieri2020tweeteval]
ade_corpus_v2-classification [gurulingappa2012development]
glue_sst [wang2018glue]
ag_news [gulli2005corpus]
paws [zhang2019paws]
glue_qqp [wang2018glue]
bbh_geometry_shapes [suzgun2022challenging]
bbh_logic_deduction_five_objects [suzgun2022challenging]
clinc
[rabinovich2022reliable]
banking77
[casanueva2020efficient]
NLI
Training: 6 datasets
Test: 4 datasets
circa [louis2020d]
scitail [khot2018scitail]
art [bhagavatula2019abductive]
glue_qnli [rajpurkar2016squad]
glue_mnli [williams2017broad]
anli [nie2019adversarial]
glue_rte
[dagan2005pascal]
glue_wnli
[levesque2012winograd]
sick [marelli2014semeval]
super_cb
[dolan2005automatically]
QA
Training: 8 datasets
Test: 6 datasets
hellaswag [zellers2019hellaswag]
cosmos_qa [huang2019cosmos]
kilt_ay2 [hoffart2011robust]
quail [rogers2020getting]
dream [sun2019dream]
sciq [welbl2017crowdsourcing]
qasc [khot2020qasc]
gpqa [rein2023gpqa]
quartz_no_knowledge
[tafjord2019quartz]
quartz_with_knowledge
[tafjord2019quartz]
openbookqa [mihaylov2018can]
ai2_arc [clark2018think]
codah [chen2019codah]
quarel [tafjord2019quarel]
Multilingual SUM
Training: 7 datasets
Test: 2 datasets
xlsum en, xlsum zh, xlsum jp,
xlsum fr, xlsum ru, xlsum es, xlsum ko [hasan-etal-2021-xl]
xlsum pt [hasan-etal-2021-xl]
xlsum sr [hasan-etal-2021-xl]
Appendix B Computational Cost Analysis
B.1 Training cost
We compare mask all targets training cost with that of mask last token in terms of complexity analysis and quantitative estimation.
B.1.1 Complexity analysis
Our proposed mask all targets strategy is significantly more efficient than the standard mask last target approach for training ICL capabilities. Mask last target needs separate sequences for 1-shot, 2-shot,…n-shot, processing, with $O(n\times n_{w})$ tokens per instance, where $n$ is the maximum number of shots and $n_{w}$ is the maximum token length of the context window. In contrast,mask all targets learns from 0 to n-shot from a single packed sequence, processing only $O(n_{w})$ tokens.
B.1.2 Quantitative estimation
For the training, we use $N=70K,n_{w}=32k,n=100$, and there are approximately 212 avg shots. In this setup, mask all targets processed around 2.2B tokens versus 237B for mask last target, which results in roughly a 100x reduction.
Table 7: Mask all targets and mask last target training efficiency comparison.
Training Strategy Theoretical analysis Practical estimation
[…]
We compare the computational cost of meta-training paradigm with mask last token in terms of complexity analysis and quantitative estimation. ManyICFT enables a "fine-tune only once" model. While the single fine-tuning uses long sequences (handled efficiently by mask-all-targets), it avoids the significant cumulative cost of separate fine-tuning for numerous downstream tasks. The number of training tokens are summarized in Table 8.
Table 8: Comparison on the number of training tokens for Meta training with task-level fine-tuning.
Meta learning (ManyICFT) Task-level fine-tuning
70K $\times$ 32K = 2.2B 4K $\times$ 8K $\times$ 1K = 32B
B.2 Inference cost
Table 9: Inference cost comparison: the few-shot, many-shot without KV cache and with KV cache inference theoretical complexity and its cost in practice. ($k$ is the number of shots for few-shot setting and $n$ is that for many-shot setting, $n_{x}$ denotes the average number of input tokens and $n_{y}$ is that of output tokens for each shot.)
Theoretical complexity
Practical costs for
32K context length
Few-shot $O(k^{2}(n_{x}+n_{y})^{2}$) $0.0025\times$
Many-shot w/o KV cache $O(n^{2}(n_{x}+n_{y})^{2})$ $1\times$
Many-shot w KV cache $O(n(n_{x}+n_{y})^{2}$ ) $0.01\times$
B.2.1 Complexity analysis
The vanilla inference costs is quadratic to the number of context length due to attention mechanism. With the help of KV-cache, the context can be cached and reduce the inference costs linear to the context length.
B.2.2 Practical costs
[…]
In conclusion, our mask all targets strategy makes the single ManyICFT fine-tuning phase highly efficient ( $100\times$ fewer tokens processed in our implementation than mast last target strategy). This achieves efficient fine-tuning than fine-tuning each individual task in task-level fine-tuning approaches.
Appendix C Experiment Results
C.1 Additional results on ablation study
The performance comparison among ManyICFT and baselines on many-shot inference for ablation study is reported in Table 10. This demonstrates the proposed ManyICFT still keeps best performance on ablation study. We also report zero-shot and five-shot results of ablation study in Table 11. These results show that zero-shot, few-shot, and autoregressive fine-tuning perform worse than the base model, indicating significant forgetting on out-of-domain tasks. In contrast, the many-shot fine-tuned model maintains strong performance when tested on unseen task types.
Table 10: Ablation study. Metrics are reported in both the best accuracy/accuracy with maximum in-context samples. Ablation are done by removing the corresponding training dataset of the test task. For example, when testing the performance on Clinc and Banking77, we remove the CLS dataset during training.
CLS NLI QA MATH SUM
Clinc BANKING77
No
fine-tuning
Zero-shot
FT
0.876/0.858 0.835/0.835 0.808/0.798 0.804/0.804 0.435/0.435 0.272/0.280
Few-shot
FT
0.955/0.955 0.860/0.860 0.818/0.812 0.808/0.808 0.446/0.442 0.291/0.291
Many-shot
0.964/0.964 0.880/0.880 0.844/0.844 0.818/0.814 0.454/0.448 0.296/0.296
Many-shot
(With all training data)
Table 11: Out-of-domain generalization results for zero-shot and five-shot evaluations. Accuracy is reported in the format of zero-shot/five-shot. This experiment assesses the generalization ability of different fine-tuning strategies across six distinct tasks without including task-specific training data
Clinc Banking NLI QA GSM8K SUM
No fine-tuning 0/0.535 0/0.387 0.446/0.744 0.465/0.653 0/0.411 0.170/0.264
Zero-shot 0/0.514 0/0.372 0.397/0.792 0.424/0.577 0/0.385 0.147/0.245
Few-shot
0/0.563 0/0.392 0.415/0.779 0.445/0.647 0/0.392 0.152/0.271
Many-shot
0/0.538 0/0.382 0.422/0.727 0.453/0.551 0/0.372 0.158/0.232
Many-shot
0/0.558 0/0.388 0.498/0.805 0.451/0.632 0/0.404 0.175/0.274
C.2 Analysis on the number of training samples
We provide a quantitative analysis for the number of training samples for baseline and proposed method. The results are shown in Table 12 and we can observe that ManyICFT only requires 1K in-context examples for CLS tasks while standard task-level fine-tuning requires 4K fine-tuning samples to achieve the best performance. Similarly for other tasks including QA, NLI, GSM8K and SUM, ManyICFT reduces the required samples significantly. ManyICFT leverages relatively few inference-time examples (hundreds to thousands) to achieve strong results, replacing the need for potentially much larger task-specific training sets used in standard fine-tuning. While ManyICFT requires an initial diverse dataset for its one-time fine-tuning, this cost is amortized across all tasks, avoiding the cumulative effort of sourcing and training on potentially thousands of samples for each new task individually.
Table 12: Comparison of the number of examples required to achieve the best performance.
CLS Banking77 NLI QA GSM8K SUM
Task-level FT 4K 4K 12K 18K 7.4K 4K
Many-shot FT 1.2K 0.6K 0.13K 0.2K 0.08K 0.02K
Relative
improvements
[…]